OpenStudy (anonymous):

If r(x) = −4x − 1 and t(x) = 2x − 1, find (r + t)(7). 1. −16 2. 16 3. 14 4. −14

OpenStudy (anonymous):

-16

OpenStudy (anonymous):

r(7) = −4*7 − 1 t(7) = 2*7 − 1 r(7) = -29 t(7) = 13 +___ - 16